OpenStudy (anonymous):

solve for x. 1/4x + 2/3 = x + 3/4 I have been trying to figure out how to do problems like these for over a week! So far, I have figured out pretty much only that I have to combine terms. Could I get some explanation on how to do this?

OpenStudy (saifoo.khan):

Do you know solving it with LCDs?

OpenStudy (saifoo.khan):

\[\frac14 x + \frac23 = x + \frac34\]For left side, the LCD = 12\[\frac{x(3) + 2(4)}{12} = x + \frac34\]

OpenStudy (anonymous):

Yes, I know that for fractions I have to do this so that I can use them. The LCD I believe is 12. but I don't know much after that.

OpenStudy (saifoo.khan):

Ok, so you get till here?

OpenStudy (saifoo.khan):

\[\frac{3x + 8}{12} = x + \frac34\]Now on right side, take LCD = 4. \[\frac{3x + 8}{12} = \frac{x(4) + 3}{4}\]

OpenStudy (saifoo.khan):

\[\frac{3x+8}{12} = \frac{4x + 3}{4}\]Now cross multiply, \[4(3x+8) = 12(4x+3)\]Can you solve now?

OpenStudy (anonymous):

Let me try

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

I have to simplify the numbers in the parenthesis, right?

OpenStudy (phi):

Khan has videos that might help http://www.khanacademy.org/math/algebra/solving-linear-equations/v/multi-step-equations

OpenStudy (phi):

I would distribute (that means multiply 7 times each term inside the parens) and distribute the 12 (multiply 12 times each term in the parens.)